What are the formulas to find the required informationUsing the expanded accounting equation, calculate and enter the answers for each question. You will need to use the answers y
Answer #1

Part 3)

The value of net income (loss) is arrived as below:

Ending Equity = Beginning Equity + Net Income - Dividends + Common Stock

Substituting values in the above equation, we get,

35,000 = 11,000 + Net Income - 42,000 + 7,400

Rearranging values, we get,

Net Income (Loss) = 35,000 - 11,000 + 42,000 - 7,400 = $58,600

_____

Part 4)

The value of common stock is determined as follows:

Ending Equity = Beginning Equity + Net Income - Dividends + Common Stock

Substituting values in the above equation, we get,

35,000 = 11,000 + 1,300 - 9,500 + Common Stock

Rearranging values, we get,

Common Stock = 35,000 - 11,000 - 1,300 + 9,500 = $32,200

_____

Part 5)

The value of dividends is calculated as below:

Ending Equity = Beginning Equity + Net Income - Dividends + Common Stock

Substituting values in the above equation, we get,

35,000 = 11,000 + 18,200 - Dividends + 15,000

Rearranging values, we get,

Dividends = 11,000 + 18,200 + 15,000 - 35,000 = $9,200

_____

Part 6)

The value of net income (loss) is derived as below:

Ending Equity = Beginning Equity + Net Income - Dividends + Common Stock

Substituting values in the above equation, we get,

35,000 = 11,000 + Net Income - 1,700 + 42,900

Rearranging values, we get,

Net Income (Loss) = 35,000 - 11,000 + 1,700 - 42,900 = -$17,200